PLEASE HELP ME WITH THIS PROBLEM: y with the exponent 3 - 2y with the exponent 2 + 3y -4 when y = 5 i do not under stand if in the part 2y with the exponent 2 do i multiply 2 * 5 or do the exponent with 5 first?: PLEASE HELP ME WITH THIS PROBLEM: y with the exponent 3 - 2y with the exponent 2 + 3y -4 when y = 5 i do not under stand if in the part 2y with the exponent 2 do i multiply 2 * 5 or do the exponent with 5 first?
Answer by checkley77(3848)

F(5)=Y^3-2Y^2+3Y-4
F(5)=5^3-2*5^2+3*5-4 [HERE YOU MUST FIRST RAISE TO THE POWER (5^2) THEN MULTIPLY
F(5)=125-2*25+15-4
F(5)=125-50+15-4
F(5)=86 ANS.